30 June 2012
D. Gonzalez Huerta
beginning of night
Dave worked on his mosalign script, taking some data, running it.
UT 20120630 luci.20120630.NNNN.fits
1x4.0 0001-2 O2DCR for readnoise
1x12.0 0003-4 MER10 for readnoise
from MOS ACQU on LUCI: X=1.118 Y=-2.026 Rot=0.484
shell RIPped on the way to yan1
From MOS ACQU on LUCI: X=1.029 Y=-2.567 Rot=0.302
from mosalign 0.158 -0.509 0.303
Found bug in mosalign - geomap input list needs to be around ~0,0
Changed mosalign to subtract off the coords of the mask center
No chance to test because another shell RIP that took 3+h to resolve
then clouds kept us closed for the rest of the night.
Optic Sieve to check IQ
Towards the end of this LUCI work, there were problems with shell (actually it had ripped 3 times during the first half of the night). Juan Carlos has been working on recovery from this last RIP for quite a while already - seems to be more serious than the usual.
telescope is at zenith, dome is closed and dark. Starting some MODS calibrations:
20120629 MODS calibrations
- grflats.cal will be 20120630 b:1-10 and r:1-15 (red stuck on 5, need to run red fitsflush)
- interesting... the exposure and readout countdown bars are grey in this remote display, not blue... other colors are there.
- grlamps.cal will be b: 11-13 and r:16-18
- prflats.cal (telescope was moved during this sequence...all blue, 14-23, are OK but only one r, 19, is)...
- mods1b.20120629.n.fits were n=2-5 and
- mods1r.20120629.n.fits where n=2-3
called it a night as it was overcast - already we had been closed for a while during the shell recovery (during which time the above mentioned MODS calibrations were done)
- 30 Jun 2012